In a container, two mole of a diatonic ideal gas is allowed to expand against 1 ATM pressure &volume change from 2 litre to 5 liter then calculate the change in internal energy

Grade:11

1 Answers

Vikas TU
7 years ago
Internal Energy that is U = FnR(T2-T1)/2
where F is the degree of freedom.
Therfore, F = 5 for diatomic gas,
U = 5*2*R(T2-T1/2 => 2.5*1*(5-2) => 15/2 = 7.5 Joules.